### v2

@@ Introduction
However, the rapid development of protein generative models and the lack of corresponding regulations lead to copyright and biosecurity concerns. First, the ease of model sharing raises copyright concerns, such as the unauthorized use of generated structures and the redistribution of pretrained models for profit, potentially compromising the interests of the original creators [7, 8, 9]. For example, the latest AlphaFold3 is only deployed on the server with the terms saying that users “must not use AlphaFold Server or its outputs to train machine learning models” or “in connection with any commercial activities”[10]. Second, the unregularized yet powerful protein generative models are vulnerable to misuse and cause bio-security/safety concerns [11, 12]. For example, protein generative models can be used to design new proteins with harmful properties, such as pathogens [13], toxins [14], or viruses [15] that can be used as bioweapons. Therefore, there is an urgent need for a reliable and efficient tool to track and audit the use of protein generative models.
@@ Methods
To save the computation costs of fine-tuning and flexibly embed watermark information in the generation process, we proposed the Watermark LoRA strategy. The computation formula for LoRA can be then written as $\Delta\mathbf{W}=\mathbf{A}\times\mathbf{E}\times\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{n\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times m}$, $\mathbf{E}\in\mathbb{R}^{r\times r}$ as the embedding matrix. Here, $r$ is the rank of the matrix ($r\ll n,m$). When $\mathbf{E}$ is an identity matrix, the computation yields the same result with the vanilla LoRA [35].
To input the watermark information into the fine-tuned model, we apply a mapper that transforms a watermark code of length $l$ into a vector of length $r$. Specifically, for the $i$-th bit of the code, we use a embedding vector $\mathbf{I}_{i}$ and $\mathbf{0}$ to represent the binary states 1 and 0, with $\mathbf{I}_{i},\mathbf{0}\in\mathbb{R}^{r}$. The mapping function $h_{i}:\{0,1\}\rightarrow\mathbb{R}^{r}$ is defined as
@@
In the generation process, when adding a watermark to the model for a given watermark code, we pass the watermark through the aforementioned mapper to obtain the embedding matrix $\mathbf{E}$. We then calculate the added model weights $\Delta\mathbf{W}$ and merge it into the original model weights by calculating $\mathbf{W}_{\text{watermarked}}=\mathbf{W}+\alpha\Delta\mathbf{W}$.
In the default setting, $\mathbf{I}_{i}$ is initialized from a standard normal distribution. However, to better maximize the differences between vectors, we propose using orthogonal initialization. Specifically, for any $\mathbf{I}_{j}$ and $\mathbf{I}_{k}$, where $j,k\in[0,l]$, we ensure that $\mathbf{I}_{j}\cdot\mathbf{I}_{k}=\mathbf{0}$. During model training, we fixed the mapper’s weights, as fine-tuning them showed no significant improvement.

### v3

@@ Introduction
However, the rapid development of protein generative models and the lack of corresponding regulations lead to copyright and biosecurity concerns. First, the ease of model sharing brings up copyright concerns, including the risk of unauthorized use of generated structures and the redistribution of pretrained models for profit, which could undermine the interests of the original creators [7, 8, 9]. For example, the latest AlphaFold3 is only deployed on the server with the terms saying that users “must not use AlphaFold Server or its outputs to train machine learning models” or “in connection with any commercial activities”[10]. Second, the unregularized yet powerful protein generative models are vulnerable to misuse and cause bio-security/safety concerns [11, 12]. For example, protein generative models can be used to design new proteins with harmful properties, such as pathogens [13], toxins [14], or viruses [15] that can be used as bioweapons. Therefore, there is an urgent need for a reliable and efficient tool to track and audit the use of protein generative models.
@@ Methods
To save the computation costs of fine-tuning and flexibly embed watermark information in the generation process, we proposed the Watermark LoRA strategy. The computation formula for LoRA can be expressed as $\Delta\mathbf{W}=\mathbf{A}\times\mathbf{E}\times\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{n\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times m}$, and $\mathbf{E}\in\mathbb{R}^{r\times r}$ serves as the embedding matrix. In this context, $r$ represents the rank of the matrix, satisfying $r\ll n,m$. When $\mathbf{E}$ is set as an identity matrix, the computation aligns with the vanilla LoRA setup [35].
To input the watermark information into the fine-tuned model, We utilize a mapping function that converts a watermark code of length $l$ into a vector of length $r$. For each $i$-th bit of the code, an embedding vector $\mathbf{I}_{i}$ represents the binary state 1, while $\mathbf{0}$ represents the binary state 0, where $\mathbf{I}_{i},\mathbf{0}\in\mathbb{R}^{r}$. This mapping function, $h_{i}:\{0,1\}\rightarrow\mathbb{R}^{r}$, is defined as
@@
During the generation process, when adding a watermark to the model for a given watermark code, we pass the watermark through the mapper as mentioned earlier to obtain the embedding matrix $\mathbf{E}$. We then calculate the added model weights $\Delta\mathbf{W}$ and incorporate it into the original model weights using the formula: $\mathbf{W}_{\text{watermarked}}=\mathbf{W}+\alpha\Delta\mathbf{W}$.
$\mathbf{I}_{i}$ can be initialized from a standard normal distribution. In FoldMark, we propose using orthogonal initialization to maximize the differences between vectors. Specifically, for any $\mathbf{I}_{j}$ and $\mathbf{I}_{k}$, where $j,k\in[0,l]$, we ensure that $\mathbf{I}_{j}\cdot\mathbf{I}_{k}=\mathbf{0}$. During model training, we fixed the mapper’s weights, as fine-tuning them showed no significant improvement.